How many grams of aluminum hydroxide are needed to react with 1.0 grams of hydrochloric acid?
1
Expert's answer
2013-02-28T11:20:25-0500
The chemical equation for thi reaction is


Al(OH)3 + 3HCl = AlCl3 + 3H2O
MW (HCl) = MW (H) + MW (Cl) = 1.0 + 35.5 = 36.5 g/mol
MW ( Al(OH)3 ) = MW (Al) + 3 x MW (O) + 3 x MW (H) = 27.0 + 3 x 16.0 + 3 x 1.0 = 78.0 g/mol

The amount of hydrocloric acid is

n (HCl) = 1.0 :36.5 = 0.027 mol

The amount of aluminium hydroxide

n ( Al(OH)3 ) = n (HCl) / 3 = 0.027 :3 = 9.0 x 10^-3 moles

The weight of& Al(OH)3 needed is

m ( Al(OH)3 ) =& n ( Al(OH)3 ) x MW ( Al(OH)3 ) = 9.0 x 10^-3 x 78.0 = 0.702 grams

Answer& : m ( Al(OH)3 ) = 0.702 grams